OpenStudy (anonymous):

An electric company charges $7.50 per month plus $0.09 for each kilowatt hour (kwh) of electricity used. Use C as the total cost and n as the number of kwh. Write an expression that gives the total cost of n kwh of electricity. C=7.50+0.09n You can find the average cost per kwh by dividing total cost by number of kwh. Write the expression that gives the average cost per kwh. C(kwh)=C/n Find the average cost per kwh for 10 kwh, 100 kwh, and 1,000 kwh. 10 kwh You would first start by finding the overall cost C=7.50+0.09(10) C=7.50+0.90 C=0.84 kwh Next plug in 0.84 for C and 10 for n.

OpenStudy (anonymous):

C(kwh)=0.84/10 C(kwh)=0.084 100 kwh C=7.50+0.09(100) C=7.50+9 C=16.50 C(kwh)=16.50/100 C(kwh)=0.165 1000 kwh C=7.50+0.09(1000) C=7.50+90 C=97.50 C(kwh)=97.50/1000 C(kwh)=0.0975 As the number of kwh increases, what happens to the average cost of the kwh? Would the average cost ever fall below $0.09? If so, identify a value that supports your answer. If not, explain how you know. As the number of kwh increases the average cost of the kwh decreases. How many kwh should be purchased for the average cost to be $0.12 per kwh?

OpenStudy (anonymous):

The average cost of the kwh will approach 0.09 as you use more electricity, never going below it because of the $7.50 monthly charge.

OpenStudy (anonymous):

I think that I did all the other problems right, but I am struggling with this class.

OpenStudy (anonymous):

First one looks wrong because the avg price per kwh can't be below $.09/kwh

OpenStudy (anonymous):

I appreciate the help so much. I think it is best if I just continue working on it tomorrow when I have a clear head and get some rest.

OpenStudy (anonymous):

Good idea

OpenStudy (anonymous):

The work looks OK except for the first calculation at 10kwh because you got C=.84 when it should've been C=8.4. Then the average price per kwh would've been $.84

OpenStudy (anonymous):

To figure out how many kwh need to be used for the average price per kwh to equal .12, set \[\frac{7.50+0.09n}{n}=.12\] Then you can solve for n .12n=7.50+0.09n .03n=7.50 n=250 kwh